OpenStudy (yacoub1993):

Bamboo plants grow rapidly. A bamboo plant is 130 inches tall. Tomorrow it will be 143 inches tall, the next day it will be 156 inches tall, and on the next day it will be 169 inches tall. Write an explicit formula to represent the height of the bamboo plant as an arithmetic sequence. How tall will the plant be in 13 days? A. A(n) = 130 + 13n; 286 inches B. A(n) = 130 + (n-1)13; 286 inches C. A(n) = 130 + (n-1)13; 299 inches D. A(n) = 130 + 13n; 299 inches <--------

OpenStudy (anonymous):

B is correct answer since general term of an arithematic series is A(n)= a+ (n-1)d; where a is the ist term of the series i.e in the question a=130,d is the common difference i.e in the question d=13 since 143-130=13 and 156-143=13 substituting in the mentioned formula the plant will be 286 inches tall

OpenStudy (yacoub1993):

ahaa your correct i miscalculated

OpenStudy (anonymous):

you got it

OpenStudy (yacoub1993):

no but the total hight should be 299 in am i correct

OpenStudy (yacoub1993):

its C or D i think

OpenStudy (yacoub1993):

let me tell you how i did it

OpenStudy (yacoub1993):

everyday 13 inches increases and for 13 days it will be 13 x 13 = 169 add 169 to 130 of the frist hieght 130 + 169 = 299

OpenStudy (anonymous):

May I join :)?

OpenStudy (yacoub1993):

yes you can

OpenStudy (yacoub1993):

@E.ali

OpenStudy (anonymous):

sorry i made a mistake since n=13 because it is asked thirteen days from today and the asnwer is d

OpenStudy (yacoub1993):

lets double check with @E.ali

OpenStudy (anonymous):

and the formula would be 130 +13*n

OpenStudy (anonymous):

@Yacoub1993 u are correct

OpenStudy (anonymous):

Sure ! @Yacoub1993 :First just find this : 

OpenStudy (anonymous):

OK ?:)

OpenStudy (yacoub1993):

? = 13

OpenStudy (anonymous):

Yep :) Then we have +13 for each days :) Now we know +13 is a number we have it at all . OK :)?

OpenStudy (yacoub1993):

ok then

OpenStudy (anonymous):

Now for 13 day that h will be 13*13 ok ?:)

OpenStudy (yacoub1993):

which is 169

OpenStudy (yacoub1993):

then

OpenStudy (anonymous):

Then we have this : The first hight =130 and we got 169 . To find next h : 169+130=299 And we said 13*13 so if get days n get 13 to n So We have : 130+13n=299 So u was right and D is right :)

OpenStudy (anonymous):

@Yacoub1993 :Thank both of u . @manojbalaji1:

OpenStudy (anonymous):

D is incorrect actually, the correct answer is A. A(n) = 130 + 13n; 286 inches